vectors b=13i+4j and c=17i+11j. what is the resultant vector in i+j form? what is the angle of the resultant vector with respect to the x-axis?
Answer #1

Resultanat vector = B + c

= (13+17) i + ( 4 +11 ) j

= 30 i + 15 j <=================Answere

angle of resultant vector = Tan inv(15/30)

= 26.56 degrees <==================Answere
Answer #2

resultant is b + c

b + c = 30 i + 15 j

angle = arctan(15/30) = 26.57 degrees